One side of a triangle is three times the shortest side. The third side is 7 feet more than the shortest side. The perimeter is 62 feet. Find all three sides.
Answer by mananth(16946)

shortest side be x
one side = 3x
third side= x+7
...
x+x+7+3x=62
5x+7=62
-7
5x+7-7=62-7
5x=55
/5
x= 55/5
x=11
3x = 33
11+7=18
..
11, 33, 18